# Learning when to rank: Estimation of partial rankings from sparse, noisy comparisons

### v2

@@ III.3 Case Study: CS Faculty Hiring
Figure 4: Partial rankings of CS departments according to faculty hiring patterns. (a) Barplot of BT strengths for the 205 PhD-granting institutions included in the dataset, ordered according to their inferred BT rankings. Colors indicate the PR membership of each institution. A legend displaying the numerical values of these PR strengths is shown in panel (b). The names of the five institutions making up the strongest PR cluster and of the first five institutions in the second-strongest PR cluster (ordered in terms of their BT rank) are shown in the figure. (b) Barplot depicting the inverse of the USN ranking score for the 153 institutions for which data was available. Colors indicate the corresponding PR membership of each institution. Some notable inconsistencies between the USN rankings of some institutions and those inferred by pairwise comparison methods are shown in the figure.
@@
Figure 5: Partial rankings of CS departments and their external metadata. (a) BT rankings as a function of USN rankings for all 153 PhD-granting institutions that had associated USN ranking. Points are colored according to their PR rank. The black dashed line corresponds to the case $R_{BT}=R_{USN}$. (b) Average BT strength of a node’s out-neighbors as a function of the node’s out-degree. Points are colored according to their PR rank. The horizontal lines represent the average BT score of the out-neighbors disaggregated by PR group. (c) Average endowment (in millions of USD) for the institutions in each partial rank. Error bars represent the standard deviations within each partial rank group. (d) Average endowment (in millions of USD) as a function of the USN rank, disaggregated by partial rankings. Error bars represent the standard deviations within each group of universities binned by USN rank (including ties). (e) Network representation of faculty hiring flows among the partial rank groups inferred by our algorithm. Each node represents a partial ranking, with labels and colors indicating the corresponding group. Node sizes are proportional to the number of institutions assigned to each partial ranking. Directed edges represent the hiring flows between groups, with edge thickness proportional to the net flow. Edge colors transition from the origin node’s color at the base to the destination node’s color at the arrowhead and the net flow value is annotated on each edge.
@@
As a case study, we apply our algorithm to a dataset of faculty hires in Computer Science (CS) departments at U.S. higher education institutions, originally compiled in clauset2015systematic for analyzing hierarchies in academic hiring using a different ranking methodology. Each node in the network represents one of the 205 PhD-granting institutions included in the study, and a directed edge $(i,j)$ is included for each individual who received their doctorate from institution $i$ and held a tenure or tenure-track faculty position at institution $j$ during the data collection period (May 2011 to March 2012). For 153 of these institutions, we have access to their U.S. News & World Report (USN) rankings, which we use as node metadata to compare against the rankings and partial rankings inferred by the two pairwise ranking methods explored in this study. The USN rankings are based on a broad set of criteria—including reputation, research output, student selectivity, and resources—and, as such, are not expected to directly reflect hiring dynamics. However, it is nonetheless useful to examine the extent to which these widely cited rankings align with those inferred from the academic job market dynamics. Additionally, we manually collected endowment data for 151 of the 153 institutions with USN rankings to explore potential correlations with inferred prestige.
@@
Fig. 4(a) shows the results of applying the BT model and our PR algorithm to the entire set of 205 PhD-granting institutions. Each bar represents an institution, ordered by its BT rank, with bar heights representing the BT strength scores. Bars are color-coded according to their PR rankings, with corresponding strengths shown in the legend in Fig. 4(b). The rankings are dominated by a small group of elite universities displaying significantly higher strengths, followed by a sharp decline in strength values further down the list. The PR and BT rankings show strong ordinal association ($\tau_{B}=0.82$; p-value $=1.2\times 10^{-53}$) with only a small number of rank violations—instances where $\text{rank}_{BT}(A)>\text{rank}_{BT}(B)$ but $\text{rank}_{PR}(A)<\text{rank}_{PR}(B)$— at the boundary between the two lowest-scoring PR groups, possibly indicating a lack of information to accurately determine the group boundaries.
@@
One first thing to notice with regards to the USN rankings is that they allow for partial rankings. For example, the four highest-scoring universities according to the USN methodology are all assigned an equal rank of 1. Fig. 4(b) reveals that institutions with equal USN ranking often belong to the same partial rankings according to the PR model. However, the PR model generally produces broader rankings, grouping institutions that span a wider range of USN ranks. Another striking feature is the significantly higher number of rank violations, although the overall ordinal association remains strong ($\tau_{B}=0.76$; p-value $=1.37\times 10^{-33}$). Some of these rank violations are particularly striking. For example, Harvard University, ranked $5^{th}$ in the BT model and belonging to the top-scoring PR group, is ranked $17^{th}$ according to the USN rankings, behind institutions including the University of Maryland, College Park ($26^{th}$ in the BT model and in the third PR group) that scored significantly lower according to the hiring network structure. Similarly, the University of Chicago is ranked $19^{th}$ by the BT model and belongs to the second highest-scoring PR group but is ranked $35^{th}$ in the USN rankings. While the U.S. News & World Report utilizes a wide array of metrics to conduct its rankings, such as peer assessment scores, graduation and retention rates, and financial resources, this result would appear to suggest that these factors may not directly align with academic job market preferences. In the case of Harvard University or the University of Chicago, one might surmise that the prestigious reputation of the institutions might be influencing the job market more than their USN rankings. The opposite effect can also be observed. For instance, Northeastern University, ranked $61^{st}$ in the USN rankings, is assigned $178^{th}$ place by the BT model and belongs to the lowest PR group despite its considerably higher USN ranking. This suggests that perhaps this program was undervalued in the job market relative to its academic credentials during the period studied.
@@
To evaluate how the BT rankings compare to the USN rankings, Fig. 5(a) plots the BT rank of each institution against its corresponding USN rank, with points color-coded based on their PR rank. The highest-ranking institutions, according to both models, tend to concentrate around the $R_{BT}=R_{USN}$ line, where $R_{BT}$ and $R_{USN}$ represent the BT and USN rankings respectively—indicating close agreement between the two ranking systems for top-tier institutions. As we move towards lower-ranked institutions, we observe an increasing divergence between the rankings. While the overall ordinal association remains strong ($\tau_{B}=0.73$, p-value $=1.07\times 10^{-38}$), the spread in rankings widens, and institutions with a given USN rank are often assigned considerably different BT ranks, with the BT model frequently assigning these institutions higher positions compared to the USN rankings. This discrepancy is particularly pronounced among institutions in the two lowest-scoring PR groups.
@@
An interesting question about the pairwise ranking methods considered in this work is whether an institution’s rank benefits more from “who you win against” than from the sheer number of “wins”. In other words, does an institution’s ranking improve more by producing a large number of doctoral graduates who secure tenure-track positions across a wide range of institutions or by producing fewer graduates who secure positions at highly ranked institutions? To address this, in Fig. 5(b), we plot the average BT strength of an institution’s out-neighbors (capturing the quality of an institution’s “wins”) as a function of its out-degree (representing the total number of “wins” an institution has). We observe a clear positive correlation (Pearson $p=0.78$; p-value $=1.32\times 10^{-43}$) with the highest ranked institutions having both a large number of “wins” (i.e. producing a large number of graduates that go on to secure tenure-track positions at other institutions), as well as out-neighbors with high BT strengths (i.e. their graduates tend to secure positions at high-ranking institutions). Perhaps more revealing is the fact that the PR groups appear to stratify according to the average BT strength of their out-neighbors, suggesting a well-defined hierarchy characterized by a predominance in horizontal mobility as opposed to vertical mobility. This observation is reinforced by analyzing the network of mobility flows between the PR groups; see Fig. 5(e). The results reveal a well-defined hierarchy with minimal upward mobility across ranks. Most mobility occurs downward within the hierarchy, a trend partly explained by the relatively small size of the top-ranked groups. Nonetheless, horizontal mobility emerges as a critical driver of hiring dynamics across all levels, particularly within the lower-ranked groups, where it becomes the dominant hiring pattern. Horizontal mobility is also particularly pronounced at the topmost level of the hierarchy, where the top five institutions comprising this elite group hire over twice as many faculty members from within their group as they do from all other PhD-granting institutions combined.
@@
Finally, we note that the ranking patterns possess a moderately strong ordinal association with the endowments of the institutions. Although there is considerable variability, higher-ranked institutions generally tend to have significantly larger endowments, as illustrated in Fig. 5 (c)-(d).
@@ IV Conclusion
Finally, while we have developed a fully Bayesian nonparametric framework to infer partial rankings, we have focused on inferring point estimates of these rankings without considering uncertainty quantification as done in other works pearce2024bayesian ; peixoto2022ordered . However, as noted in Sec. II.3, the Bayesian nature of our model lends itself naturally to the exploration of the full posterior, which can be explored via the development of suitable inference algorithms—such as Expectation-Maximization (EM) or Markov Chain Monte Carlo (MCMC) methods—to enable posterior sampling or approximation, thereby allowing for principled uncertainty quantification in inferred rankings.
@@ Acknowledgments
The authors thank Max Jerdee for useful discussions about the datasets, and Michael Pearce and Elena Erosheva for helpful suggestions to improve the initial manuscript. The authors acknowledge funding support from the HKU-100 Start Up Fund and an HKU Urban Systems Institute Fellowship Grant.
@@ Appendix A Minimization of $f(R)$
We begin proving that the second term in Eq. 39 is maximized when all groups are of equal size such that $n_{r}=n=N/R~{}~{}\forall r$.
@@
$\binom{N}{n_{1}\ldots n_{R}}$ $=N\log N-N-\sum_{r=1}^{R}(n_{r}\log n_{r}-n_{r})$ (42)
$=N\log N-\sum_{r=1}^{R}n_{r}\log n_{r},$ (43)
@@
To maximize Eq. 40, we then have to solve the following constrained optimization problem
@@
($\mathcal{L}(\{n_{r}\},\lambda)=-\sum_{r=1}^{R}n_{r}\log n_{r}+\lambda\left(\sum_{r=1}^{R}n_{r}-N\right),$ (46)
@@
Given Proposition 1 and in view of the monotonicity of the logarithm, it follows that the second term on the right-hand side of Eq. 39 is maximized when all groups are of equal size.
@@
From Proposition 1 and by virtue of the monotonicity of the logarithm, it follows that Eq. 39 is upper bounded by
@@
where we have once more made use of Stirling’s approximation. The first term on the right-hand side of Eq. 52 is a concave function of $R$, reaching its minimum value of zero at both $R=1$ and $R=N$. The second term, however, increases monotonically with $R$, starting from a minimum of zero when $R=1$. Since these two terms combine additively in Eq. 52, their sum must attain its global minimum of zero at $R=1$, where both terms simultaneously achieve their lowest values. Finally, since the binomial and multinomial coefficients are always greater than or equal to one, it follows that Eq 39 is also lower bounded by zero, meaning it must attain a global minimum of zero at $R=1$. ∎
@@ Appendix B Time Complexity
Appendix C Tau Recovery

### v3

@@ III.3 Case Study: CS Faculty Hiring
Figure 5: Partial rankings of CS departments according to faculty hiring patterns. (a) Barplot of BT strengths ($y$-axis) for the 205 PhD-granting institutions included in the dataset, ordered according to their inferred BT rankings ($x$-axis). Colors indicate the PR membership of each institution. A legend displaying the numerical values of these PR strengths is shown in panel (b). The names of the five institutions making up the strongest PR cluster and of the first five institutions in the second-strongest PR cluster (ordered in terms of their BT rank) are shown in the figure. (b) Barplot depicting the PR groups (colored as before) with respect to the USN ranking ($x$-axis) instead of BT rank, for the 153 institutions for which USN data was available. Some notable inconsistencies between the USN rankings of some institutions and those inferred by pairwise comparison methods are shown in the figure. The inverse of USN ranking was plotted along the $y$-axis to provide an analogue to the BT score for these pre-determined rankings for easier visualization.
@@
Figure 6: Partial rankings of CS departments and their external metadata. (a) BT rankings as a function of USN rankings for all 153 PhD-granting institutions that had an associated USN ranking. Points are colored according to their PR rank. The black dashed line corresponds to the case $R_{BT}=R_{USN}$. (b) Average BT strength of a node’s out-neighbors as a function of the node’s out-degree. Points are colored according to their PR rank. The horizontal lines represent the average BT score of the out-neighbors disaggregated by PR group. (c) Average endowment (in millions of USD) for the institutions in each partial rank. Error bars represent the standard deviations within each partial rank group. (d) (Top panel) Average endowment (in millions of USD) as a function of the USN rank, disaggregated by partial rankings. Error bars represent the standard deviations within each group of universities binned by USN rank (including ties). (Bottom panel) Strip plot of the the number of institutions within each (USN rank, PR rank) group, i.e. the number of institutions belonging to each bar. (e) Network representation of faculty hiring flows among the partial rank groups inferred by our algorithm. Each node represents a partial ranking, with labels and colors indicating the corresponding group. Node sizes are proportional to the number of institutions assigned to each partial ranking. Directed edges represent the hiring flows between groups, with edge thickness proportional to the net flow. Edge colors transition from the origin node’s color at the base to the destination node’s color at the arrowhead and the net flow value is annotated on each edge.
@@
As a case study, we apply our algorithm to a dataset of faculty hires in Computer Science (CS) departments at U.S. higher education institutions, originally compiled in clauset2015systematic for analyzing hierarchies in academic hiring using a different ranking methodology. Each node in the network represents one of the 205 PhD-granting institutions included in the study, and a directed edge $(i,j)$ is included for each individual who received their doctorate from institution $i$ and held a tenure or tenure-track faculty position at institution $j$ during the data collection period (May 2011 to March 2012). For 153 of these institutions, we have access to their U.S. News & World Report (USN) rankings, which we use as node metadata to compare against the rankings and partial rankings inferred by the two pairwise ranking methods explored in this study.
The USN rankings are constructed from a broad set of indicators—including reputation surveys, research output, student selectivity, and institutional resources—and therefore reflect a metrics-based assessment of institutional prestige. In contrast, the rankings inferred from hiring data capture a more behavioral notion of prestige, one that emerges from the collective decisions of faculty members and hiring committees. Comparing these two types of rankings allows us to identify not only their areas of agreement but also systematic differences between them, revealing how the hierarchies implied by hiring dynamics may diverge from those defined by traditional metrics. These differences highlight the distinction between perceived prestige, i.e., how academics act upon their professional judgments when hiring or accepting positions, and evaluated prestige, i.e., how institutions are ranked based on quantifiable performance indicators such as publications, grants, and resources. Additionally, we manually collected endowment data for 151 of the 153 institutions with USN rankings to explore potential correlations with inferred prestige.
@@
Fig. 5(a) shows the results of applying the BT model and our PR algorithm to the entire set of 205 PhD-granting institutions. Each bar represents an institution, ordered by its BT rank, with bar heights representing the BT strength scores. Bars are color-coded according to their PR rankings, with corresponding strengths shown in the legend in Fig. 5(b). The rankings are dominated by a small group of elite universities displaying significantly higher strengths, followed by a sharp decline in strength values further down the list. The PR and BT rankings show strong ordinal association ($\tau_{B}=0.82$; p-value $1.2\times 10^{-53}$) with only a small number of rank violations—instances where $\text{rank}_{BT}(A)>\text{rank}_{BT}(B)$ but $\text{rank}_{PR}(A)<\text{rank}_{PR}(B)$— at the boundary between the two lowest-scoring PR groups, possibly indicating a lack of information to accurately determine the group boundaries.
@@
One first thing to notice with regards to the USN rankings is that they allow for partial rankings. For example, the four highest-scoring universities according to the USN methodology are all assigned an equal rank of 1. Fig. 5(b) reveals that institutions with equal USN ranking often belong to the same partial rankings according to the PR model. However, the PR model generally produces broader rankings, grouping institutions that span a wider range of USN ranks, suggesting that hiring-based prestige hierarchies are consistent with, but less granular than, those defined by metrics-based rankings, reflecting a structure of academic reputation that operates at the level of broad institutional tiers rather than fine rank distinctions. Another striking feature is the significantly higher number of rank violations, although the overall ordinal association remains strong ($\tau_{B}=0.76$; p-value $1.37\times 10^{-33}$). Unlike in the BT case, where rank violations likely result from limited information, the presence of rank violations across the entire ranking spectrum suggests that they reflect genuine differences between evaluated prestige—as measured by metrics-based rankings—and perceived prestige—as revealed through hiring dynamics. Indeed, some of these rank violations are particularly striking. For example, Harvard University, ranked $5^{th}$ in the BT model and belonging to the top-scoring PR group, is ranked $17^{th}$ according to the USN rankings, behind institutions including the University of Maryland, College Park ($26^{th}$ in the BT model and in the third PR group) that scored significantly lower according to the hiring network structure. Similarly, the University of Chicago is ranked $19^{th}$ by the BT model and belongs to the second-highest-scoring PR group, but is ranked $35^{th}$ in the USN rankings, suggesting that these factors may not directly align with academic job market preferences. In the case of Harvard University or the University of Chicago, one might surmise that the prestigious reputation of the institutions might be influencing the job market more than their USN rankings. The opposite effect can also be observed. For instance, Northeastern University, ranked $61^{st}$ in the USN rankings, is assigned $178^{th}$ place by the BT model and belongs to the lowest PR group despite its considerably higher USN ranking. This suggests that perhaps this program was undervalued in the job market relative to its academic credentials during the period studied.
@@
To evaluate how the BT rankings compare to the USN rankings, Fig. 6(a) plots the BT rank of each institution against its corresponding USN rank, with points color-coded based on their PR rank. The highest-ranking institutions, according to both models, tend to concentrate around the $R_{BT}=R_{USN}$ line, where $R_{BT}$ and $R_{USN}$ represent the BT and USN rankings respectively—indicating close agreement between the two ranking systems for top-tier institutions. As we move towards lower-ranked institutions, we observe an increasing divergence between the rankings. While the overall ordinal association remains strong ($\tau_{B}=0.73$, p-value $1.07\times 10^{-38}$), the spread in rankings widens, and institutions with a given USN rank are often assigned considerably different BT ranks, with the BT model frequently assigning these institutions higher positions compared to the USN rankings. This discrepancy is particularly pronounced among institutions in the two lowest-scoring PR groups. In general, these results suggest that, outside the top-ranked institutions, subjective perceptions of prestige can significantly shape professional mobility, elevating or diminishing institutions relative to their metric-based ranking. Likewise, the observation that institutions with similar BT ranks span a broad range of USN ranks indicates that perceived prestige, as expressed through hiring and faculty mobility, does not always correspond directly to quantitative measures of institutional performance.
@@
An interesting question about the pairwise ranking methods considered in this work is whether an institution’s rank benefits more from “who you win against” than from the sheer number of “wins”. In other words, does an institution’s ranking improve more by producing a large number of doctoral graduates who secure tenure-track positions across a wide range of institutions or by producing fewer graduates who secure positions at highly ranked institutions? To address this, in Fig. 6(b), we plot the average BT strength of an institution’s out-neighbors (capturing the quality of an institution’s “wins”) as a function of its out-degree (representing the total number of “wins” an institution has). We observe a clear positive correlation (Pearson $p=0.78$; p-value $1.32\times 10^{-43}$) with the highest ranked institutions having both a large number of “wins” (i.e. producing a large number of graduates that go on to secure tenure-track positions at other institutions), as well as out-neighbors with high BT strengths (i.e. their graduates tend to secure positions at high-ranking institutions). Perhaps more revealing is the fact that the PR groups appear to stratify according to the average BT strength of their out-neighbors, suggesting a well-defined hierarchy characterized by a predominance in horizontal mobility as opposed to vertical mobility. This observation is reinforced by analyzing the network of mobility flows between the PR groups; see Fig. 6(e). The results reveal a well-defined hierarchy with minimal upward mobility across ranks. Most mobility occurs downward within the hierarchy, a trend partly explained by the relatively small size of the top-ranked groups. Nonetheless, horizontal mobility emerges as a critical driver of hiring dynamics across all levels, particularly within the lower-ranked groups, where it becomes the dominant hiring pattern. Horizontal mobility is also particularly pronounced at the topmost level of the hierarchy, where the top five institutions comprising this elite group hire over twice as many faculty members from within their group as they do from all other PhD-granting institutions combined.
@@
Finally, we note that the ranking patterns possess a moderately strong ordinal association with the endowments of the institutions. Although there is considerable variability, higher-ranked institutions generally tend to have significantly larger endowments, as illustrated in Fig. 6 (c)-(d). The high variability can be partly attributed to differences in the number of institutions contributing to each (USN rank, PR rank) group—i.e., the number of universities represented by each bar. Group sizes vary considerably across the data set, although most bars consist of relatively few institutions (1–25 universities per bar; median = 2, mode = 1). As a result, standard deviations are generally uninformative for most bars, with variability estimates becoming more reliable only for larger groups.
@@ IV Conclusion
Finally, while we have developed a fully Bayesian nonparametric framework to infer partial rankings, we have focused on inferring point estimates of these rankings without considering uncertainty quantification as done in other works pearce2024bayesian ; peixoto2022ordered . However, as noted in Sec. II.3, the Bayesian nature of our model lends itself naturally to the exploration of the full posterior, which can be explored via the development of suitable inference algorithms—such as Expectation-Maximization (EM) or MCMC methods—to enable posterior sampling or approximation, thereby allowing for principled uncertainty quantification in inferred rankings.
@@ Acknowledgments
The authors thank Max Jerdee for useful discussions about the datasets, and Michael Pearce and Elena Erosheva for helpful suggestions to improve the initial manuscript.
Data accessibility
All data used in the study are open source and freely available on the Internet. We have provided links and references in the text and electronic supplementary material for all the data used in the study.
Conflict of interest declaration
We declare we have no competing interests.
Funding
The authors acknowledge funding support from the HKU-100 Start Up Fund and an HKU Urban Systems Institute Fellowship Grant.
@@ Appendix A Minimization of $f(R)$
We begin proving that the second term in Eq. (40) is maximized when all groups are of equal size such that $n_{r}=n=N/R~~\forall r$.
@@
$\binom{N}{n_{1}\ldots n_{R}}$ $=\exp\left\{N\log N-N-\sum_{r=1}^{R}(n_{r}\log n_{r}-n_{r})\right\}$ (43)
$=\exp\left\{N\log N-\sum_{r=1}^{R}n_{r}\log n_{r}\right\},$ (44)
@@
To maximize Eq. (41), we then have to solve the following constrained optimization problem
@@
$\mathcal{L}(\{n_{r}\},\lambda)=-\sum_{r=1}^{R}n_{r}\log n_{r}+\lambda\left(\sum_{r=1}^{R}n_{r}-N\right),$ (47)
@@
Given Proposition 1 and in view of the monotonicity of the logarithm, it follows that the second term on the right-hand side of Eq. (40) is maximized when all groups are of equal size.
@@
From Proposition 1 and by virtue of the monotonicity of the logarithm, it follows that Eq. (40) is upper bounded by
@@
where we have once more made use of Stirling’s approximation. The first term on the right-hand side of Eq. (53) is a concave function of $R$, reaching its minimum value of zero at both $R=1$ and $R=N$. The second term, however, increases monotonically with $R$, starting from a minimum of zero when $R=1$. Since these two terms combine additively in Eq. (53), their sum must attain its global minimum of zero at $R=1$, where both terms simultaneously achieve their lowest values. Finally, since the binomial and multinomial coefficients are always greater than or equal to one, it follows that Eq (40) is also lower bounded by zero, meaning it must attain a global minimum of zero at $R=1$. ∎
@@ Appendix B Time Complexity
Appendix C Increasing $R$
Figure 8: Partial rankings in synthetic datasets with $N=50$ and $R=15$. (a) Heatmap of the number of rankings $R$ inferred by the partial rankings model, as a function of the parameters $(\sigma,\langle m\rangle)$ of the synthetic model in Sec. III.1 when $N=50$ and $R=15$. (b) Heatmap of the log posterior odds ratio (Eq. (38)) between the BT and the partial rankings model across the simulations. Positive values indicate a preference for the partial rankings model, and negative values a preference for the BT model. (c) The top panel displays the average number of rankings inferred by the best-performing model (blue) and our partial rankings model (red) as a function of the network density $\langle m\rangle$ of the network of matches for the slice at $\sigma=3.0$. The bottom panel displays the log posterior odds ratio (normalized per node) between the two models. Negative values of this difference indicate a preference for the BT model and positive values a preference for the partial rankings model. All results were obtained by averaging over 2020 different simulations from the synthetic network model of Sec. C with $N=50$ and $R=15$, and error bars indicate 22 standard errors in the mean.
Figure 9: Contributions to the inference objective. Behavior of the log posterior $\mathcal{L}$ and the various contributions to $\mathcal{L}$$C(R)$, $g(r)$, and f(r,s)f(r,s)— as a function of the inferred number of rankings $R$ for different values of $\langle m\rangle$ and planted rankings $R^{*}$. The black vertical dashed lines indicate the global minimum of $\mathcal{L}$.
We analyze the recovery performance of our algorithm as the number of partial rankings increases. Fig. 8 shows the behavior of the model for a network with $N=50$ nodes and $R=15$ planted rankings for different values of $\langle m\rangle$ and $\sigma$
By increasing the number of planted rankings $R$, we observe that the number of matches required to successfully recover the correct ranking structure also increases, especially at low noise levels ($\sigma$). This is expected, as distinguishing a larger number of groups demands more evidence in the form of pairwise comparisons.
At low $\sigma$, the algorithm continues to accurately recover the planted rankings across a wide range of conditions. However, the model still yields a less parsimonious description of the data compared to the BT model as observed in Fig. 2. This is due to the inconsistency between the extensive number of rank groups and the prior, and can in principle be addressed with a weaker prior, e.g. one that is uniform over rank partitions rather than the number of ranks.
Increasing the score separation at high $R$ values pushes the algorithm to find a larger number of ranks than expected. This arises from the behavior of the log-likelihood in the high-$\langle m\rangle$ regime as the number of planted rankings $R$ increases. Figure 9 illustrates the log posterior $\mathcal{L}$, along with the contributions of its individual components—namely $C(R)$, $g(r)$, and f(r,s)f(r,s)—as functions of the number of inferred rankings $R$, for varying values of $\langle m\rangle$ and planted rankings $R^{*}$. Figures 9(a) and (b) show the behavior of the algorithm at low and high values of $\langle m\rangle$, respectively, for $R^{*}=5$ planted rankings with equally spaced scores.
At low $\langle m\rangle$, the three terms contribute to produce a global minimum of the log posterior at $R=R^{*}=5$, thereby correctly identifying the true number of rankings. In contrast, in the high-$\langle m\rangle$ regime, the contributions from $C(R)$ and $g(r)$ diminish, and the log-posterior becomes dominated by the likelihood term $f(r,s)$. This marks the transition to a high-data regime where the influence of the prior effectively “washes out”, at which point our PR model behaves similarly to the BT model—which is prone to overfitting. Specifically, the likelihood term $f(r,s)$ decreases monotonically with increasing $R$, favoring more complex models.
However, as seen in Fig. 9(b), this decrease is sufficiently slow that the prior still regularizes the model adequately, and the global minimum remains at $R=5$. But this balance deteriorates as $R^{*}$ increases. While the behavior at low $\langle m\rangle$ remains qualitatively unchanged (Fig. 9(c)), in the high-$\langle m\rangle$ regime the monotonic decrease of $f(r,s)$ with $R$ becomes steeper. As a result, the regularizing influence of the prior weakens further, reducing its ability to constrain model complexity and allowing the algorithm to overfit.
Appendix D Tau Recovery